OpenStudy (anonymous):

Solve: 3(x + 1) + 6 = 33 10 8 8.7 −8

OpenStudy (anonymous):

You first use distributive law on the bracket, resolve the constants on LHS. Now shift the constant to RHS and solve.

OpenStudy (anonymous):

c ?

OpenStudy (anonymous):

3(x + 1) + 6 = 33 3x + 3 + 6 = 33 3x + 9 = 33 3x = 33 - 9 3x = 24 x = 24/3 x = 8